OpenStudy (anonymous):

What are the amplitude, period, phase shift, and midline of f(x) = −3 sin(4x − ð) + 2? Amplitude: 3; period: pi over 2; phase shift: x = pi over 4; midline: y = 2 Amplitude: −3; period: pi over 2; phase shift: x = pi over 2; midline: y = 2 Amplitude: 2; period: pi over 4; phase shift: x = pi over 4; midline: y = −3 Amplitude: 2; period: pi over 4; phase shift: x = pi over 2; midline: y = −3

OpenStudy (anonymous):

A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ • (cos Θ + 1). If Θ = 60°, what is the area of the opening? 1.2 in2 7.9 in2 16.3 in2 20.8 in2

OpenStudy (anonymous):

What cosine function represents an amplitude of 2, a period of 2π, a horizontal shift of π, and a vertical shift of −1? f(x) = −1 cos πx + 2 f(x) = −1 cos (x − π) + 2 f(x) = 2 cos (x − π) − 1 f(x) = 2 cos πx − 1

OpenStudy (anonymous):

The temperature of a chemical reaction ranges between –10°C and 50°C. The temperature is at its lowest point when t = 0 and completes one cycle over a six-hour period. What is a sine function that would model this reaction? f(t) = 30 sin pi over 3x + 20 f(t) = 60 sin pi over 3t + 30 f(t) = 30 sin 6t + 20 f(t) = 60 sin 6t + 30 @ganeshie8

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

What'd you get for the first one? Amplitude is always positive, that should help cancel out one of the options.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I just need help in these five questions! @ganeshie8

OpenStudy (anonymous):

Please @ganeshie8

OpenStudy (anonymous):

Thought you guys might wanna know about the first one: